# Can Video Large Multimodal Models Think Like Doubters—or Double-Down: A Study on Defeasible Video Entailment

### v1

@@ Abstract
Video Large Multimodal Models (VLMMs) have made impressive strides in understanding video content, but they often struggle with abstract and adaptive reasoning—the ability to revise their interpretations when new information emerges. In reality, conclusions are rarely set in stone; additional context can strengthen or weaken an initial inference. To address this, we introduce Defeasible Video Entailment (DVidE), a new task that challenges models to think like doubters, constantly updating their reasoning based on evolving evidence. In DVidE, given a video premise and a textual hypothesis, models must determine whether a new update strengthens or weakens the hypothesis (classification version) or generate a coherent update that modifies the entailment relationship (generation version). For solving the classification task, we propose the Chain of Counterfactual Thought framework, utilizing counterfactual reasoning, ASR-enhanced video content, and rationale refinement to reduce inference bias. For the generation task, we develop a framework that combines ASR output with a Large Language Model (LLM) to produce coherent, contextually relevant updates aligned with the intended strengthener or weakener goals. Additionally, we introduce a novel benchmark dataset, with strengthener/weakener annotations and an LLM-based evaluation metric specifically designed for assessing generative performance. Experimental results demonstrate significant improvements, highlighting our proposed method in enhancing dynamic reasoning capabilities of VLMMs.
@@ 1 Introduction
Video understanding has been a long-standing challenge in computer vision, encompassing tasks such as action recognition (Materzynska et al., 2019; Ghadiyaram et al., 2019; Wang et al., 2019a), video captioning (Lei et al., 2020; Wang et al., 2019b; Han et al., 2023), and video retrieval (Hendricks et al., 2017; Gao et al., 2017). Now, with the development of Video Large Multimodal Models (VLMMs) (Maaz et al., 2024a, b), video understanding tasks have become the key benchmark to test the visual ability of VLMMs. While existing VLMMs perform well in extracting perceptual details such as objects, activities, and temporal cues (Li et al., 2024; Chen et al., 2023b), they often struggle with higher-level abstract reasoning needed to fully interpret complex scenes. Different from video understanding tasks, the video entailment (Liu et al., 2020) task requires a model to assess whether a video premise supports or contradicts a textual hypothesis, typically assuming a fixed, definitive relationship between them. For example, in a sports video, if the hypothesis states, “The player scored a goal,” the task requires to determine whether the video confirms or refutes this claim.
Figure 1: An example of defeasibility in video entailment.
@@
However, real-world reasoning is typically defeasible, meaning conclusions can be strengthened or weakened by new information. A video clip may initially support one interpretation, but additional cues, such as a line of dialogue from Automatic Speech Recognition (ASR) or a subtle shift in visual context, may alter that conclusion. As illustrated in Figure 1, initially, Rachel’s frustration seems to stem from Gavin’s inappropriate behavior of staring at women. However, if new information reveals that he was actually on the phone with his mother discussing women’s body shapes, this strengthens the inference. Conversely, if he was just pretending Rachel was his secretary, this could suggest that her frustration stems from being misrepresented rather than from his behavior of staring at women.
@@
To address these challenges, we introduce Defeasible Video Entailment (DVidE), an extension of video entailment. Specifically, DVidE requires not only determining whether the initial relationship between a video and a hypothesis is entailment or contradiction, but also continuously revising that inference, either strengthening or weakening it, as new updates arrive. We explore this problem through two interrelated sub-tasks: (1) Classification Task: Given a video premise, a textual hypothesis, and a textual update, predict whether the update functions as a strengthener (increasing the plausibility of the hypothesis) or a weakener (making the hypothesis less likely). (2) Generation Task: Given a video premise, a hypothesis, and a goal (Weakener/Strengthener), generate a contextually grounded (Weakener/Strengthener) that meaningfully alters the original entailment relationship.
@@
In our preliminary experiments, we found that the performance of existing VLMMs was unsatisfactory for DVidE due to several key limitations: (1) Inability to Fully Capture Audio Information: Current VLMMs frequently struggle with audio-related tasks, failing to accurately comprehend and interpret audio semantics in the video. (2) Weak Visual Reasoning Ability: Biases in VLMM training can lead to limitations in visual reasoning, resulting in generated reasoning text that lacks diversity, particularly in out-of-distribution scenarios (Wang et al., 2024). (3) Poor Language Reasoning Ability: Vision-based fine-tuning can sometimes degrade an LLM’s original generative and reasoning capabilities (Yan et al., 2024; Chai et al., 2024).
@@
To address these limitations, we propose a Chain of Counterfactual Thought method to explore the diverse logical semantic relationships among the update, premise, and hypothesis, thereby reducing susceptibility to inherent inference bias in the classification task. Our approach consists of three key components. First, Counterfactual Thought-induced Rationale Generation explores diverse reasoning processes through counterfactual reasoning, enabling VLMMs to consider multiple perspectives and improve inference diversity. Second, ASR-enhanced Video Content Integration leverages an external ASR tool to supplement VLMMs’ perception of audio information, ensuring a more comprehensive understanding of video content. Third, Rationale Refinement and Selection utilizes the strong language reasoning capabilities of LLMs to refine the generated rationales, systematically removing inference errors introduced by VLMMs, and thereby improving final classification accuracy. Additionally, we incorporate ASR-enhanced Video Content Integration to refine video descriptions before prompting an LLM to refine the update based on the hypothesis, goal, and generated update, enhancing VLMMs’ generative capabilities for our generation task.
@@
To evaluate VLMMs, we employ annotators to construct the Defeasible Video Entailment benchmark. In addition to traditional generation metrics, we introduce a novel Large Language Model (LLM)-based evaluation metric specifically designed for our generation task. Experimental results demonstrate that both our classification method and generation method achieve significant improvements across multiple metrics. On the classification task, our approach achieves an accuracy improvement of 29.7% over the Video-ChatGPT model (Maaz et al., 2024a); On the generation task, our method increases the proportion of correctly generated updates from 7.50% to 60,63% on VideoLLaMA (Cheng et al., 2024).
@@
Our main contributions can be summarized as follows: (1) We introduce a benchmark for defeasible video entailment, augmenting existing video entailment datasets with strengthener/weakener annotations and extending video-language reasoning. (2) We develop a novel Chain of Counterfactual Thought classification framework that leverages counterfactual reasoning to explore the diverse entailment semantic relation. (3) We propose a generation framework that integrates the external ASR tool and LLM to produce coherent and contextually relevant updates while ensuring they align with the goal (Weakener/Strengthener). (4) We propose a novel LLM-based evaluation method for our generation task. Our experimental results demonstrate that our proposed methods significantly improve both classification and generation performance on DVidE.
@@ 2 Related Work
Defeasible Inference Tasks. Defeasible inference extends traditional entailment (Bos and Markert, 2005; Dagan et al., 2005; MacCartney and Manning, 2009; Bowman et al., 2015) by considering how new information can make the previous hypothesis more or less likely. Although standard Natural Language Inference (NLI) datasets, like SNLI (Bowman et al., 2015) and MNLI (Williams et al., 2018), focus on fixed premise-hypothesis relationships, recent work has explored defeasibility, where a piece of new information can strengthen or weaken the hypothesis. For example, the $\delta$-NLI dataset (Rudinger et al., 2020) specifically studies changes in entailment by providing additional textual information that makes the hypothesis more or less plausible. Similarly, $\delta$-CAUSAL (Cui et al., 2024) extends defeasibility to causal reasoning, evaluating how new facts influence inference dynamics. While these datasets introduce text-based defeasible inference, they do not incorporate multimodal reasoning. SNLI-VE (Xie et al., 2019) extends textual entailment to the visual domain by determining whether a hypothesis is entailed or contradicted by an image. The DVE dataset (Zhang et al., 2024) introduces defeasibility based on the SNLI-VE, allowing new textual information to alter image-hypothesis relationships. However, DVE remains limited to single-frame contexts, lacking the temporal visual and auditory information for models to judge whether the new information strengthen or weaken the hypothesis. Our work extends defeasible inference to video, providing models with richer temporal and multimodal context to improve understanding.
Video Understanding Tasks. Video understanding covers a range of tasks, such as activity recognition, event localization, and video-language reasoning. Activity recognition (Soomro et al., 2012; Kuehne et al., 2011; Piergiovanni and Ryoo, 2018; Roitberg et al., 2021) focuses on identifying human actions within videos. Event localization (Heilbron et al., 2015; Idrees et al., 2017; Tian et al., 2018; Geng et al., 2023) aims to detect the temporal boundaries of specific events, which is essential for video summarization and highlight detection. In addition, Video-language reasoning (Lei et al., 2018; Xiao et al., 2021; Xu et al., 2016; Miech et al., 2019) involves integrating textual and visual information to answer questions, generate descriptions, or make logical inferences about video content. While these tasks contribute significantly to video understanding, they primarily focus on static interpretation—classifying actions, detecting events, or retrieving relevant textual descriptions—without modeling how new evidence alters existing inferences. In real-world settings, reasoning is inherently defeasible, requiring conclusions to adapt as new evidence emerges. However, current benchmarks rarely capture such dynamic inference, limiting their relevance for tasks that require continual reasoning updates.
@@ 3 Problem Formulation
Following the definition outlined in (Rudinger et al., 2020), we define our Defeasible Video Entailment task as follows
Given a video clip premise $V$, a hypothesis $H$ is defeasible if there exists an update ${U}$ (consistent with ${V}$) such that a human would find ${H}$ less likely to be true after learning ${U}$. Specifically, an update ${U}$ is called a weakener if given a premise ${V}$ and hypothesis ${H}$, a human would most likely find ${H}$ less likely to be true after learning ${U}$; if they find ${H}$ more likely to be true, then we call ${U}$ a strengthener. Notably, the strengthener or weakener can not contradict the premise.
Classification Task. The goal of the classification task is to find a classification model $\mathcal{M}_{c}$ which predicts the update type based on the premise $V$, hypothesis $H$, and update $U$, formulated as $\hat{L}=\mathcal{M}_{c}(V,H,U)$, where $\hat{L}\in\{w,s,h\}$ denotes the predicted update type. Specifically, $\hat{L}=s$ (strengthener) is assigned if $U$ makes the hypothesis $H$ more likely to be true given the video clip $V$, and $\hat{L}=w$ (weakener) is assigned if $U$ makes the hypothesis $H$ less likely to be true given $V$.
Generation Task. The task is to build a model $\mathcal{M}_{g}$ that generates an textual update $\hat{U}$ based on the input video clip $V$, hypothesis $H$, and goal $G\in\{w,s\}$ (i.e. weakener or strengthened), formulated as $\hat{U}=\mathcal{M}_{g}(V,H,G)$, where $\hat{U}$ is the generated update.
@@ 4.1 Data Source
To build our Defeasible Video Entailment benchmark, we start with VIOLIN (Liu et al., 2020), a large-scale video-and-language inference dataset offering labeled video-hypothesis pairs for entailment classification. We then augment these pairs with textual updates that either strengthen or weaken the hypothesis, thus introducing the concept of defeasibility into video-based reasoning.
Specifically, VIOLIN consists of 95,322 (video premise, text hypothesis) pairs from 15,887 video clips, spanning over 582 hours of video. The videos are sourced from two domains: 1) popular TV shows and 2) movie clips from YouTube channels. Therefore, the topics of the video source are very comprehensive. For each video clip, annotators wrote multiple hypotheses that can be categorized as real statements or fake statements. A hypothesis is considered real if it is supported by the video and fake if it contradicts the video’s content. To incorporate defeasibility into the video domain, we extend VIOLIN by introducing annotations and adjusting video clips to construct (video premise, text update, text hypothesis) triplets. The detailed process is given in the next section.
4.2 Data Annotation
To create neutral (video premise, hypothesis) pairs, we remove the sub-clip evidence in the video premise that directly supports or contradicts the hypothesis, thereby removing explicit evidence. Specifically, if the sub-clip evidence appears at the beginning or end of the video, the remaining portion is retained as the premise. If the evidence is located in the middle, we select the longer of the two remaining segments to maintain coherence. The hypotheses from VIOLIN are used as our hypotheses, while the premises are derived from the edited video segments with the evidence removed. The edited video clips ensure that the premise no longer inherently supports or contradicts the statement, making it neutral.
To simplify the updated annotation process, we directly ask annotators to write textual updates based on the removed sub-clip, allowing our neutral (video premise, hypothesis) pairs to become real or fake statements. This is because the textual information serves as a substitute for the removed sub-clips semantics, providing direct evidence for entailment relation inference. For real statements, annotators are asked to manually write updates (strengtheners) based on the removed sub-clips to make the hypothesis more likely. Similarly, for fake statements, we write updates (weakeners) based on the removed sub-clips to make the hypothesis less likely. This annotation method ensures that the updates are grounded in the removed portion of the video, while remaining consistent with the original premise.
To enhance the reliability of our data, each sample in our benchmark is annotated by three independent annotators. Specifically, the first annotator generates an update based on our neutral premise-hypothesis pair and the corresponding removed sub-clip. The second annotator then reviews the update to ensure it meets the required criteria; if any discrepancies are found, they provide corrections to align it with the standards. Finally, a third annotator conducts a final review, making further modifications if necessary to ensure accuracy and consistency.
4.3 Statistics and Comparison
Table 1: Statistics of the DvidE dataset.
@@ 4.3 Statistics and Comparison
Table 2: Comparison of DVidE and related datasets.
@@
   Charades (Sigurdsson et al., 2016)
   InfiniBench (Ataallah et al., 2024)
   Kinetics (Kay et al., 2017)
@@
   SNLI (Bowman et al., 2015)
   MNLI (Williams et al., 2018)
   $\delta$-NLI (Rudinger et al., 2020)
   $\delta$-CAUSAL (Cui et al., 2024)
   SNLI-VE(Xie et al., 2019)
   DVE (Zhang et al., 2024)

### v2

@@ Abstract
Video Large Multimodal Models (VLMMs) have achieved impressive progress in video understanding, yet they remain limited in adaptive reasoning—the ability to revise conclusions when new evidence emerges. In reality, conclusions are rarely set in stone; additional context can strengthen or weaken an initial inference. To capture this capability, we introduce Defeasible Video Entailment (DVidE), a new task that challenges models to think like doubters, requiring them to dynamically update entailment judgments as evidence evolves. In DVidE, given a video premise and a textual hypothesis, the models must determine whether a new update strengthens or weakens the hypothesis (classification task) or generate a coherent update that modifies the entailment relationship (generation task). For solving the classification task, we propose the Chain of Counterfactual Thought (CoCT) framework, utilizing counterfactual reasoning, ASR-enhanced video content, and rationale refinement to reduce inference bias. For the generation task, we develop a framework that combines ASR output with a Large Language Model (LLM) to produce coherent, contextually relevant updates aligned with the intended strengthener or weakener goals. We also introduce a benchmark dataset with human annotations and an LLM-based evaluation metric for assessing generative performance. Experimental results show substantial improvements across multiple VLMM baselines, demonstrating that CoCT significantly enhances both classification and generation performance and strengthens the dynamic reasoning capabilities of video-language models.
Can Video Large Multimodal Models Think Like Doubters—or Double-Down: A Study on Defeasible Video Entailment
Yue Zhang   Jilei Sun   Yunhui Guo   Vibhav Gogate Department of Computer Science The University of Texas at Dallas, Richardson, TX, USA {yue.zhang, jilei.sun, yunhui.guo, vibhav.gogate}@utdallas.edu
@@ 1 Introduction
Video understanding has long been a central challenge in artificial intelligence, encompassing tasks such as action recognition Materzynska et al. (2019); Ghadiyaram et al. (2019), video captioning Wang et al. (2019); Han et al. (2023), and video retrieval Hendricks et al. (2017); Gao et al. (2017). Recent progress in Video Large Multimodal Models (VLMMs) Maaz et al. (2024a, b) has led to substantial advances in these tasks, making them primary benchmarks for assessing multi-modal visual understanding. Yet, despite their strong ability to capture perceptual information such as objects, activities, and temporal cues Li et al. (2024); Chen et al. (2023b), they often struggle with higher-level abstract reasoning needed to fully interpret complex or ambiguous scenes. Unlike conventional video understanding tasks, video entailment Liu et al. (2020) evaluates whether a video premise supports or contradicts a textual hypothesis, assuming a fixed and deterministic relation between them. For example, in a sports video, if the hypothesis states, “The player scored a goal”, the model must decide whether the visual evidence confirms or refutes that statement.
Figure 1: A defeasibility example in video entailment.
@@
Real-world reasoning, however, is typically defeasible, meaning conclusions can be strengthened or weakened by new information. A video clip may initially support one interpretation, but additional cues, such as dialogue obtained via Automatic Speech Recognition (ASR) or a subtle shift in visual context, may alter that conclusion. As illustrated in Figure 1, initially, Rachel’s frustration seems to stem from Gavin’s inappropriate behavior of staring at women. However, if new information reveals that he was actually on the phone with his mother discussing women’s body shapes, this strengthens the inference. Conversely, if he was just pretending Rachel was his secretary, this could suggest that her frustration stems from being misrepresented rather than from his behavior of staring at women.
@@
To capture this dynamic reasoning, we introduce Defeasible Video Entailment (DVidE), an extension of video entailment. Specifically, DVidE requires not only determining whether the initial relationship between a video and a hypothesis is entailment or contradiction, but also continuously revising that inference, either strengthening or weakening it, as new updates arrive. We explore this problem through two interrelated sub-tasks: (1) Classification, where the model predicts whether a given textual update strengthens or weakens the plausibility of a hypothesis given a video premise; and (2) Generation, where the model produces a contextually grounded update that meaningfully alters the entailment relation in the desired direction (strengthener or weakener).
@@
Our initial experiments revealed that existing VLMMs perform poorly on DVidE due to three key limitations: (1) Inability to Fully Capture Audio Information: Current VLMMs frequently struggle with audio-related tasks, failing to accurately comprehend and interpret audio semantics in the video. (2) Weak Visual Reasoning Ability: Biases in VLMM training can lead to limitations in visual reasoning, resulting in generated reasoning text that lacks diversity, particularly in out-of-distribution scenarios Wang et al. (2024). (3) Poor Language Reasoning Ability: Vision-based fine-tuning can sometimes degrade an LLM’s original generative and reasoning capabilities Yan et al. (2024); Chai et al. (2024).
@@
To overcome these limitations, we propose the Chain of Counterfactual Thought (CoCT) framework, which explores diverse logical relationships among the video premise, hypothesis, and update to mitigate inference bias in classification. CoCT consists of three key components. First, Counterfactual Thought-induced Rationale Generation explores diverse reasoning processes through counterfactual reasoning, enabling VLMMs to consider multiple perspectives and improve inference diversity. Second, ASR-enhanced Video Content Integration leverages an external ASR tool to supplement VLMMs’ perception of audio information, ensuring a more comprehensive understanding of video content. Third, Rationale Refinement and Selection utilizes the strong language reasoning capabilities of LLMs to refine the generated rationales, systematically removing inference errors introduced by VLMMs, and thereby improving final classification accuracy. Additionally, we incorporate ASR-enhanced Video Content Integration to refine video descriptions before prompting an LLM to refine the update based on the hypothesis, goal, and generated update, enhancing VLMMs’ generative capabilities for our generation task.
@@
To evaluate VLMMs, we employ annotators to construct the Defeasible Video Entailment benchmark. In addition to traditional generation metrics, we introduce a novel Large Language Model (LLM)-based evaluation metric specifically designed for our generation task. Experimental results demonstrate that both our classification method and generation method achieve significant improvements across multiple metrics. On the classification task, our approach achieves an accuracy improvement of 29.7% over the Video-ChatGPT model Maaz et al. (2024a); On the generation task, our method increases the proportion of correctly generated updates from 7.50% to 60,63% on VideoLLaMA Cheng et al. (2024).
@@
Our contributions can be summarized as: 1) We introduce a benchmark for defeasible video entailment, augmenting existing video entailment datasets with strengthener/weakener annotations and extending video-language reasoning. 2) We develop a novel Chain of Counterfactual Thought classification framework that leverages counterfactual reasoning to explore the diverse entailment semantic relations. 3) We propose a generation framework that integrates the external ASR tool and LLM to produce coherent and contextually relevant updates while ensuring they align with the goal (Weakener/Strengthener). 4) We propose an LLM-based evaluation method for our generation task. Our experimental results demonstrate that our proposed methods significantly improve both classification and generation performance on DVidE.
@@ 2 Related Work
Defeasible Inference Tasks. Defeasible inference extends traditional entailment Bos and Markert (2005); Dagan et al. (2005); MacCartney and Manning (2009); Bowman et al. (2015) by considering how new information can make the previous hypothesis more or less likely. Although standard Natural Language Inference (NLI) datasets, like SNLI Bowman et al. (2015) and MNLI Williams et al. (2018), focus on fixed premise-hypothesis relationships, recent work has explored defeasibility, where a piece of new information can strengthen or weaken the hypothesis. For example, the $\delta$-NLI dataset Rudinger et al. (2020) specifically studies changes in entailment by providing additional textual information that makes the hypothesis more or less plausible. Similarly, $\delta$-CAUSAL Cui et al. (2024) extends defeasibility to causal reasoning, evaluating how new facts influence inference dynamics. While these datasets introduce text-based defeasible inference, they do not incorporate multimodal reasoning. SNLI-VE Xie et al. (2019) extends textual entailment to the visual domain by determining whether a hypothesis is entailed or contradicted by an image. The DVE dataset Zhang et al. (2025) introduces defeasibility based on the SNLI-VE, allowing new textual information to alter image-hypothesis relationships. However, DVE remains limited to single-frame contexts, lacking the temporal visual and auditory information for models to judge whether the new information strengthen or weaken the hypothesis. Our work extends defeasible inference to video, providing models with richer temporal and multimodal context to improve understanding.
Video Understanding Tasks. Video understanding covers a range of tasks, such as activity recognition, event localization, and video-language reasoning. Activity recognition Soomro et al. (2012); Kuehne et al. (2011); Piergiovanni and Ryoo (2018); Roitberg et al. (2021) focuses on identifying human actions within videos. Event localization Heilbron et al. (2015); Idrees et al. (2017); Tian et al. (2018); Geng et al. (2023) aims to detect the temporal boundaries of specific events, which is essential for video summarization and highlight detection. In addition, Video-language reasoning Lei et al. (2018); Xiao et al. (2021); Xu et al. (2016); Miech et al. (2019) involves integrating textual and visual information to answer questions, generate descriptions, or make logical inferences about video content. While these tasks contribute significantly to video understanding, they primarily focus on static interpretation—classifying actions, detecting events, or retrieving relevant textual descriptions—without modeling how new evidence alters existing inferences. In real-world settings, reasoning is inherently defeasible, requiring conclusions to adapt as new evidence emerges. However, current benchmarks rarely capture such dynamic inference, limiting their relevance for tasks that require continual reasoning updates.
@@ 3 Problem Formulation
Following the definition outlined in Rudinger et al. (2020), we define our Defeasible Video Entailment task as follows
Given a video clip premise $V$, a hypothesis $H$ is defeasible if there exists an update U{U} (consistent with V{V}) such that a human would find H{H} less likely to be true after learning U{U}. Specifically, an update U{U} is called a weakener if given a premise V{V} and hypothesis H{H}, a human would most likely find H{H} less likely to be true after learning U{U}; if they find H{H} more likely to be true, then we call U{U} a strengthener. Importantly, the update must remain consistent with the video premise: it may modify belief in the hypothesis but cannot contradict the observed evidence.
[…]
Generation Task. The objective is to train a model $\mathcal{M}_{g}$ that generates a textual update $\hat{U}$ based on the input video clip $V$, hypothesis $H$, and goal $G\in\{w,s\}$ (i.e. weakener or strengthener), formulated as $\hat{U}=\mathcal{M}_{g}(V,H,G)$.
@@ 4.1 Data Source
To build our Defeasible Video Entailment benchmark, we start with VIOLIN Liu et al. (2020), a large-scale video-and-language inference dataset offering labeled video-hypothesis pairs for entailment classification. We then augment these pairs with textual updates that either strengthen or weaken the hypothesis, thus introducing the concept of defeasibility into video-based reasoning.
@@
Table 1: Statistics of the DVidE dataset.
Specifically, VIOLIN consists of 95,322 (video premise, text hypothesis) pairs from 15,887 video clips, spanning over 582 hours of video. The videos are sourced from two domains: 1) popular TV shows and 2) movie clips from YouTube channels. Therefore, the topics of the video source are very comprehensive. For each video clip, the annotators wrote multiple hypotheses that can be categorized as real statements or fake statements. A hypothesis is considered real if it is supported by the video or fake if it contradicts the video’s content. To incorporate defeasibility into the video domain, we extend VIOLIN by introducing annotations and adjusting video clips to construct (video premise, text update, text hypothesis) triplets. The detailed process is given in the next section.
4.2 Data Annotation
To create neutral (video premise, hypothesis) pairs, we remove the sub-clip evidence in the video premise that directly supports or contradicts the hypothesis, thereby removing explicit evidence. Specifically, if the sub-clip evidence appears at the beginning or end of the video, the remaining portion is retained as the premise. If the evidence is located in the middle, we select the longer of the two remaining segments to maintain coherence. The hypotheses from VIOLIN are used as our hypotheses, while the premises are derived from the edited video segments with the evidence removed. The edited video clips ensure that the premise neither inherently supports nor contradicts the statement, making it neutral.
@@ 4.2 Data Annotation
Charades Sigurdsson et al. (2016)
InfiniBench Ataallah et al. (2024)
Kinetics Kay et al. (2017)
@@
SNLI Bowman et al. (2015)
MNLI Williams et al. (2018)
$\delta$-NLI Rudinger et al. (2020)
$\delta$-CAUSAL Cui et al. (2024)
SNLI-VE Xie et al. (2019)
DVE Zhang et al. (2025)
@@
Table 2: Comparison of DVidE with related datasets.
To simplify the updated annotation process, we directly ask annotators to write textual updates based on the removed sub-clip, allowing our neutral (video premise, hypothesis) pairs to become real or fake statements. This is because the textual information serves as a substitute for the removed sub-clips semantics, providing direct evidence for entailment relation inference. For real statements, annotators are asked to manually write updates (strengtheners) based on the removed sub-clips to make the hypothesis more likely. Similarly, for false hypotheses, we construct weakening updates based on the removed sub-clips to make the hypothesis less likely to hold. This annotation strategy ensures that each update is grounded in the omitted portion of the video while remaining consistent with the original premise.
To enhance the reliability of our data, each sample in our benchmark is annotated by three independent annotators. Specifically, the first annotator generates an update based on our neutral premise-hypothesis pair and the corresponding removed sub-clip. A second annotator reviews each update to verify that it satisfies the required criteria and provides corrections if discrepancies are found. Finally, a third annotator performs a final pass, making additional adjustments as needed to ensure overall accuracy and consistency.
4.3 Statistics and Comparison